PUTSOMEMORE:

What is the GCF of 44j5k4 and 121j2k6?

barstho:

So, do you know that a GCF is?

TheSmartOne:

She obviously doesn't which is why she's asking So we have two concepts: factors and multiples The factors of a number, let's say 12 for example, would be 1, 2, 3, 4, 6, and 12 Because 1*12 = 12 2*6 = 12 3*4 = 12 Now multiples on the other hand would be the number multiplied by any number So 2 has many multiples 2, 3, 6, 8, 10, and goes on Because 2*1 = 2 2*2 = 4 And so on until infinity So like you can say that 400 is a multiple of 2 because 2*200 = 400

TheSmartOne:

So we first want to break apart our numbers 44 and 121 Can you list all the numbers that if you multiply them, that you'll get that number? And then we try to find the greatest common factor in between them. Here's an example 90 and 100 90 has the factors 1, 3, 9, 10, 30 and 90 100 has the factors 1, 2, 4, 5, 10, 10, 20, 25, 50, 100 What's the largest number in common between them? 10 So 10 is our greatest common factor And that basically tells us that we can factor out 10 from both of those numbers? So how does this help us? Let's say you had something like 90x + 100y And you want to simplify it, then you can because you know the greatest common factor 10(9x +10y)

TheSmartOne:

So can you first list all the factors for 44 and 121? And then try to find the greatest common factor. I'll be here to help you out along the way

TheSmartOne:

And then for the variables, that is the j and k We just take the greatest common factor So for example if you have \(j^2k\) and \(jk^2\) They both have one j and one k in common so the greatest common factor is jk This is a very simple example, but I can elaborate more once you find the GCF of 44 and 121

kittybasil:

@TheSmartOne the person may know what a GCF is but not how to get one, js 😂

TheSmartOne:

... what is your point exactly? I'm not sure if you've read my entire response